OpenStudy (anonymous):

O\left (\exp \left ( \left (\tfrac{64n}{9} \log(2) \right )^{\frac{1}{3}} \left ( \log(n\log(2)) \right )^{\frac{2}{3}} \right) \right )

OpenStudy (anonymous):

OK lets see

OpenStudy (anonymous):

what the heck is this?

OpenStudy (anonymous):

\[O\left (\exp \left ( \left (\tfrac{64n}{9} \log(2) \right )^{\frac{1}{3}} \left ( \log(n\log(2)) \right )^{\frac{2}{3}} \right) \right )\]

OpenStudy (anonymous):

this I weird

OpenStudy (anonymous):

someone knows how to answer that?